cos(180-theta)
cos(270-theta)
cos(360-theta)
of sum and difference of two angle.So can u please give me proof for the above questions.

Answer by Alan3354(69443)

cos(180-theta)
cos(270-theta)
cos(360-theta)
of sum and difference of two angle.
----------
cos(A-B) = cos(A)*cos(B) + sin(A)*sin(B)
------------
cos(180-t) = cos(180)*cos(t) + sin(180)*sin(t)
= -1*cos(t) + 0
= -cos(t)
=================
Do the others the same way.